OpenStudy (anonymous):

HELP PLEASE! A sound wave is modeled with the equation y =1/4 cos 2pi/3 theta a. Find the period. Explain your method.

OpenStudy (anonymous):

I know the amplitude is 1/4 but I don't know how to find the period.

OpenStudy (anonymous):

Would the period be 3?

OpenStudy (sidsiddhartha):

if there is afunction \[y=acosbx\] then period of this function is \[T=\frac{ 2\pi }{ b}\] so here the function is y =1/4 cos 2pi/3 theta so period=\[\large\frac{ 2\pi }{ \frac{ 2\pi }{ 3}}=\large3\]

OpenStudy (sidsiddhartha):

do u get it? @amypperson

OpenStudy (anonymous):

Yes thank you soo much :)

OpenStudy (sidsiddhartha):

^_^